Circular No. 7453

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 2000 W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova in the Large Magellanic Cloud on two unfiltered Kodak Technical Pan films obtained using a 0.2-m Schmidt camera on July 12.4 UT. The object was not present on a film he had taken of the region on 1999 Mar. 8. On July 13.0, using a CCD at the Schmidt’s Newtonian focus, he measured the object’s position as α = 5^{h}25^{m}01^{s}.60, δ = −70^{o}14′17′′.3 (equinox 2000.0). On July 13.4, a low-dispersion objective-prism CCD spectrogram, also using the Schmidt, showed the object to have intense Hα emission, with a strength > 100 times that of the surrounding continuum. Hβ emission was also clearly present. Observed magnitudes: July 12.4098 UT, 11.2 (Liller); 12.4129, 11.2 (Liller); 12.98, V = 11.45 (Liller, comparison GSC 9166.0178); 13.469, m_{v} = 11.8 (R. Stubbings, Drouin, Vic., Australia, comparison GSC 9166.0673; VY Hyi sequence suggests m_{v} = 12.2).

SUPERNOVA 2000cu IN ESO 525-G004 M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope of an apparent supernova on unfiltered images taken on July 12.3 (mag ∼ 16.2) and 13.3 UT (mag ∼ 16.1). The object is located at α = 19^{h}24^{m}20^{s}.95, δ = −25^{o}59′04′′.1 (equinox 2000.0), which is 12′′.1 east and 1′′.5 north of the nucleus of ESO 525-G004. An unfiltered image of the field taken on July 5.3 showed nothing at the position of the object (limiting mag ∼ 19.0).

COMETS C/2000 M6, 2000 M7, 2000 M8 D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports observations of three more Kreutz sungrazing comets detected with the SOHO/LASCO coronagraphs (C2 alone for the first, both C3 and C2 for the others). Detections were by M. Meyer, by H. Luethen and Meyer and by Biesecker, respectively. Discovery positions follow. Complete measurements (by D. Hammer, reduction by B. G. Marsden) and orbital elements are given on MPEC 2000-N31.

2000 UT  α_{2000}  δ_{2000}  Comet
June 25.479  6^{h}12.^{m}9  +21^{o}46′  C/2000 M6
26.471  6 14.9  +20 26  C/2000 M7
27.779  6 20.8  +21 18  C/2000 M8
2000 July 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden